# DAWN-FM: Data-Aware and Noise-Informed Flow Matching for Solving Inverse Problems

### v2

@@ 2.3 A data-aware and noise-informed velocity estimator
To this end, notice that we train a velocity function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ that takes in two arguments. In the context of a specific inverse problem, we have additional information for training the network, the measured data ${\bf b}$ and the noise level $\sigma$ in ${\bf b}$. Note that it is also possible to estimate the noise level directly from the data as proposed in Tenorio et al. [2011]. The data ${\bf b}$ can be used to point toward ${\bf x}_{1}$ even at time $t=0$, where ${\bf x}_{t}$ contains no information on ${\bf x}_{1}$ and can therefore improve the estimation of velocity. Moreover, the information about noise level $\sigma$ in the measured data during training also makes the estimator more robust to noise during inference. We thus propose to use the data and noise when estimating ${\bf s}_{{\boldsymbol{\theta}}}$. To this end, we use a transformation $f$ of data ${\bf b}$ (to be discussed next), and let
@@
$\widehat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}\mathcal{L}_{1}{({\boldsymbol{\theta}})}={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\Big{[}\left\|{\bf s}_{\theta}({\bf x}_{t},f({\bf A}{\bf x}_{1}+\sigma{\bf z}),t,\sigma)+{\bf x}_{0}-{\bf x}_{1}\right\|^{2}\Big{]},$ (9)
@@
As we show in our numerical experiments, having $\sigma$ as an input to the network plays an important role, generating an inversion methodology that is robust to different noise levels.
@@
This approach was used in Mardani et al. [2018], Adler and Öktem [2017], and as shown in Section 2.5, can be successful for the transformation of the data into the image space. Other possible approaches can include fast estimation techniques for ${\bf x}$ given ${\bf b}$ such as the conjugate gradient least squares method [Hansen, 1997]. For the experiments presented here, we found that using the adjoint ${\bf A}^{\top}$ of the forward problem was sufficient.
@@
Example 2.1
The Duathlon problem: We consider the duathlon problem where one records the total time it takes to perform a duathlon (bike and run). Given the total time, the goal is to recover the time it takes to perform each individual segment. The problem is clearly under-determined as one has to recover two number given a single data point. Let ${\bf x}=[x_{1},x_{2}]$ be the vector, where $x_{1}$ and $x_{2}$ represent the time it takes to finish the bike and run segments, respectively. The data is simply
$b=x_{1}+x_{2}+\epsilon$
@@
Now, in order to solve the inverse problem, we train a larger network that includes the data and approximately solves the optimization problem in Equation 9. Given the data ${\bf b}$, we now integrate the ODE Equation 10 to obtain a solution. The result of this integration is presented in Figure 2 (right). We observe that not only did the process identify the correct lobe of the distribution, it also sampled many solutions around it. This enables us in obtaining not just a single but rather an ensemble of plausible solutions that can aid in exploring uncertainty in the result.
@@
Figure 2: Generation of the distribution of two Gaussians by FM without and with data.
Figure 3: Schematic of the training process for the FM model for solving inverse problem, where the forward model is given by ${\bf A}$. This figure specifically represents the deblurring inverse problem. The network with parameters ${\boldsymbol{\theta}}$ can either have noise embedding (DAWN-FM) or not (DAW-FM). The two loss terms $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the error in prediction of velocity and misfit, respectively. In this figure, the transformation $f$ for generating the data embedding was chosen as $f={\bf A}^{\top}$.
2.4 Training the FM model for inverse problems
@@
${\bf b}_{{\boldsymbol{\theta}}}={\bf A}{\bf x}_{t}+(1-t){\bf A}{\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},f({\bf b}),t,\sigma)$ (12)
If the velocity is estimated exactly, then ${\bf b}_{{\boldsymbol{\theta}}}={\bf b}$. Therefore, a natural loss for the recovery is the comparison of the recovered data to the given data. This implies that the velocity function should honor the data ${\bf b}$, as well as the original distribution ${\bf x}_{1}$. We thus introduce the misfit loss term,
@@
$\widehat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}\Big{\{}{{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\|{\bf s}_{\theta}({\bf x}_{t},f({\bf b}),t,\sigma)-{\bf v}\|^{2}}+\alpha{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\|{\bf b}_{{\boldsymbol{\theta}}}-{\bf b}\|^{2}\Big{\}},$ (14)
@@
2.5 Architectures for data-aware and noise-informed velocity estimators
Our goal is to train a velocity estimator ${\bf s}_{\theta}({\bf x}_{t},{\bf b},t,\sigma)$ and then compare it to the velocity ${\bf v}$. Our estimator is based on a UNet [Ronneberger et al., 2015] with particular embeddings for ${\bf b},t$, and $\sigma$.
The embedding of time and noise is straightforward. Note that both $t$ and $\sigma$ are scalars. For their embedding, we use the method presented in Croitoru et al. [2023], which involves creating learnable embedding and adding them to the feature maps at each level of the network.
A key component in making FM perform well for inverse problems is the careful design of a network that incorporates the data within the training process. The resulting network can be thought of as a likelihood-free estimator [Thomas et al., 2022], that estimates the velocity vector ${\bf v}$, given the input vector ${\bf x}_{t}$, the time $t$ and the noise level $\sigma$. As previously explained, we do not integrate ${\bf b}$ directly, but rather use $f({\bf b})={\bf A}^{\top}{\bf b}$ and embed this vector in the network.
@@ 2.5 Architectures for data-aware and noise-informed velocity estimators
be an encoder network that is parameterized by ${\boldsymbol{\eta}}$. The encoder can be pre-trained or trained as part of the network, providing flexibility in how it is integrated into the UNet. The encoded data provides additional context that is critical for accurate predictions. Formally, the UNet estimates the velocity as,
$\widehat{\bf v}={\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},E_{{\boldsymbol{\eta}}}({\bf A}^{\top}{\bf b}),t,\sigma)$ (16)
@@
2.6 Inference and uncertainty estimation
Figure 4: Computing uncertainty in the solutions obtained by FM for some example images from MNIST (left), STL10 (middle) and CIFAR10 (right) for the deblurring task. The reconstruction (posterior mean) is computed by averaging over runs from 32 randomly initialized ${\bf x}_{0}$. The uncertainty in the posterior mean is computed by evaluating its standard deviation for the predictions from 32 runs.
Our method is specifically aimed at highly ill-posed inverse problems, where regularization is often necessary to arrive at a stable solution. These problems often do not have a unique solution, and slight changes in input data can lead to large variations in the solution. Due to its stochastic nature, our method allows for the realization of multiple solutions since one can start with many random initial points ${\bf x}_{0}$ and evolve to multiple versions ${\bf x}_{1}$ of the solutions. This property can be used to generate even better estimates and for the estimation of uncertainty in the recovered images. To this end, assume that the ODE is solved $M$ times starting each time at a different initial condition ${\bf x}_{0}^{(j)}$ and let
@@ 2.6 Inference and uncertainty estimation
${\bar{\bf x}}_{1}={\frac{1}{M}}\sum_{j=1}^{M}{\bf x}_{1}^{(j)},\quad{\sigma}_{{\bar{\bf x}}_{1}}={\frac{1}{M}}\sum_{j=1}^{M}\|{\bf x}_{1}^{(j)}-{\bar{\bf x}}_{1}\|^{2}$ (18)
as the mean and standard deviation of the estimated solutions. In particular, ${\bar{\bf x}}_{1}$ approximates the posterior mean and ${\sigma}_{{\bar{\bf x}}_{1}}$ estimates its standard deviation. It is well known that, for many problems the posterior mean can have a lower risk compared to other estimators (e.g. the Maximum A-Posteriori estimator). For an elaborate discussion on the properties of such estimators, see Tenorio et al. [2011], Kaipio and Somersalo [2004], Calvetti and Somersalo [2005]. While such estimators are typically avoided due to computational complexity, the computational framework presented here allows us to compute them relatively easily.
@@
3 Numerical Experiments

### v3

@@ 2.3. A data-aware and noise-informed velocity estimator
To this end, notice that we train a velocity function ${\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},t)$ that takes in two arguments. In the context of a specific inverse problem, we have additional information for training the network, the measured data $\bf b$} and the noise level $\sigma$ in $\bf b$}. Note that it is also possible to estimate the noise level directly from the data as proposed in [Tenorio2011]. The data $\bf b$} can be used to point toward ${\bf x}_{1}$ even at time $t=0$, where ${\bf x}_{t}$ contains no information on ${\bf x}_{1}$ and can therefore improve the estimation of velocity. Moreover, the information about noise level $\sigma$ in the measured data during training also makes the estimator more robust to noise during inference. We thus propose to use the data and noise when estimating ${\bf s}_{{\boldsymbol{\theta}}}$. To this end, we use a transformation $f$ of data $\bf b$} (to be discussed next), and let
@@
$\hat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}\mathcal{L}_{1}{({\boldsymbol{\theta}})}={\rm arg}\min_{{\boldsymbol{\theta}}}{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\Big[\left\|{\bf s}_{\theta}({\bf x}_{t},f({\bf A}{\bf x}_{1}+\sigma{\bf z}),t,\sigma)+{\bf x}_{0}-{\bf x}_{1}\right\|^{2}\Big],$ (9)
@@
As we show in our numerical experiments, having $\sigma$ as an input to the network plays an important role, generating an inversion methodology that is robust to different noise levels. While our method assumes that the noise standard deviation $\sigma$ from the data is known, this is not a strict limitation. In practice, $\sigma$ can often be estimated using classical image-adaptive heuristics such residual-based [liu2006noise], wavelet-domain estimators [pimpalkhute2021digital], or inferred by analyzing the eigenvalue distribution of patch covariance matrices [chen2015efficient].
@@
This approach was used in [mardani2018neural, adler2017solving], and as shown in Section 2.5, can be successful for the transformation of the data into the image space. Other possible approaches can include fast estimation techniques for $\bf x$} given $\bf b$} such as the conjugate gradient least squares method [hansen]. For the experiments presented here, we found that using the adjoint $\bf A$$^{\top}$ of the forward problem was sufficient.
@@
Example 2.1.
The Duathlon problem: We consider the duathlon problem where one records the total time it takes to perform a duathlon (bike and run). Given the total time, the goal is to recover the time it takes to perform each individual segment. The problem is clearly under-determined as one has to recover two number given a single data point. Let ${\bf x}=[x_{1},x_{2}]$ be the vector, where $x_{1}$ and $x_{2}$ represent the time it takes to finish the bike and run segments, respectively. The data is simply,
$b=x_{1}+x_{2}+\epsilon.$
@@
Now, in order to solve the inverse problem, we train a larger network that includes the data and approximately solves the optimization problem in Equation 9. Given the data $b$, we now integrate the ODE Equation 10 to obtain a solution. The result of this integration is presented in Figure 2 (right). We observe that not only did the process identify the correct lobe of the distribution, it also sampled many solutions around it. This enables us in obtaining not just a single but rather an ensemble of plausible solutions that can aid in exploring uncertainty in the result.
@@
Figure 2. Illustration of the Duathlon problem using flow matching (FM). Each ${\bf x}=[x_{1},x_{2}]$ corresponds to bike and run times drawn from a prior distribution consisting of two Gaussians (shown in black). The red points denote observed measurements obtained through the forward model, $b=x_{1}+x_{2}+\epsilon$, and the blue curves show generated trajectories under FM. Left: FM generation without conditioning on data recovers the full mixture distribution, sampling across both Gaussian lobes. Right: FM generation with conditioning on measured data correctly identifies the appropriate lobe consistent with the observations and produces an ensemble of plausible solutions around it, thereby capturing the posterior variability of the inverse problem.
Figure 3. Schematic of the training process for the FM model for solving inverse problem, where the forward model is given by $\bf A$}. This figure specifically represents the deblurring inverse problem. The network with parameters $\boldsymbol{\theta}$} can may or may not incorporate the noise embedding (see Section 4 for details). The two loss terms $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ represent the error in prediction of velocity and misfit, respectively. In this figure, the transformation $f$ for generating the data embedding was chosen as $f={\bf A}^{\top}$.
2.4. Training the FM model for inverse problems
@@
${\bf b}_{{\boldsymbol{\theta}}}={\bf A}{\bf x}_{t}+(1-t){\bf A}{\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},f({\bf b}),t,\sigma).$ (12)
If the velocity is estimated exactly, then ${\bf b}_{{\boldsymbol{\theta}}}={\bf b}$.
Note that if the velocity field ${\bf s}_{\boldsymbol{\theta}}$ perfectly recovers the true conditional drift along the training trajectories and the measurement noise is zero (i.e., ${\bf b}={\bf A}{\bf x}_{1}$), then by construction, the reconstruction $\hat{{\bf x}}_{1}={\bf x}_{t}+(1-t){\bf s}_{\boldsymbol{\theta}}(\cdot)$ satisfies ${\bf A}\hat{{\bf x}}_{1}={\bf b}$, implying ${\bf b}_{\theta}={\bf b}$. However, this identity holds strictly in the noiseless case. In the presence of measurement noise or when ${\bf s}_{\boldsymbol{\theta}}$ is only an approximation of the true velocity field, ${\bf b}_{\boldsymbol{\theta}}$ becomes a random variable that does not necessarily equal $\bf b$} for each sample. In this case, we can enforce data consistency in expectation by minimizing the discrepancy using $\|{\bf A}\hat{{\bf x}}_{1}-{\bf b}\|^{2}$ through the $\ell_{2}$ loss. This ensures that the predicted measurements align with observed data in a statistical sense, even though exact sample-wise agreement is not guaranteed.
Therefore, a natural loss for the recovery is the comparison of the recovered data to the given data, meaning the the velocity function should honor the data $\bf b$}, as well as the original distribution ${\bf x}_{1}$. We thus introduce the misfit loss term,
@@
$\hat{\boldsymbol{\theta}}={\rm arg}\min_{{\boldsymbol{\theta}}}\Big\{{{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\|{\bf s}_{\theta}({\bf x}_{t},f({\bf b}),t,\sigma)-{\bf v}\|^{2}}+\alpha{\mathbb{E}}_{{\bf x}_{0},{\bf x}_{1},\sigma,t}\|{\bf b}_{{\boldsymbol{\theta}}}-{\bf b}\|^{2}\Big\},$ (14)
@@
It is important to note that while $\mathcal{L}_{1}$ is data-aware through its inputs ($f({\bf b})$), it does not by itself enforce that the image implied by the predicted velocity reproduces the measured data under the forward model. This is the main reason for introducing $\mathcal{L}_{2}$, which penalizes discrepancy between the observed data $\bf b$} and the data synthesized from the denoised, one-step estimate, $\hat{{\bf x}}_{1}$ passed through the forward operator, i.e., ${\bf b}_{\theta}={\bf A}\hat{{\bf x}}_{1}$. In effect, $\mathcal{L}_{2}$ is a physics/data-consistency term that anchors the learned velocities so their implied reconstructions match the actual measurements, which is especially important in the early times $t\approx 0$ (when ${\bf x}_{t}$ carries little information about ${\bf x}_{1}$) and at higher noise levels. Practically, $\mathcal{L}_{2}$ behaves like a consistency regularizer on the velocity field (not a prior on the image), i.e., it does not constraint the null-space components of the forward model $\bf A$}, but it suppresses solutions that fit $\mathcal{L}_{1}$ while violating the forward model in the range of $\bf A$}. This complementary interplay between $\mathcal{L}_{1}$ (for amortized probability transport) and $\mathcal{L}_{2}$ (for explicit measured data fidelity) proved useful in our experiments.
Figure 4. Example images from MNIST (left) and STL10 (right) datasets showing the recovery of deblurred images from the blurred (noisy) data (top panel) at different levels of noise (0%, 1%, 5%, 10%, 20%). Our methods DAW-FM  (middle panel) incorporates the embedding for the blurred data within the network, while DAWN-FM (bottom panel) incorporates embedding for both data and noise levels. DAWN-FM is superior to DAW-FM at deblurred image recovery, especially at higher levels of noise (see Section 4 for detailed implementation details for DAW-FM and DAWN-FM architectures.)
2.5. Architectures for data-aware and noise-informed velocity estimators
Our goal is to train a velocity estimator ${\bf s}_{\theta}({\bf x}_{t},{\bf b},t,\sigma)$ and then compare it to the velocity $\bf v$}. Our estimator is based on a UNet [UNET2015] with particular embeddings for ${\bf b},t$, and $\sigma$.
The embedding of time and noise is straightforward. Note that both $t$ and $\sigma$ are scalars. For their embedding, we use the method presented in [croitoru2023diffusion], which involves creating learnable embedding and adding them to the feature maps at each level of the network.
A key component in making FM perform well for inverse problems is the careful design of a network that incorporates the data within the training process. The resulting network can be thought of as a likelihood-free estimator [thomas2022likelihood], that estimates the velocity vector $\bf v$}, given the input vector ${\bf x}_{t}$, the time $t$ and the noise level $\sigma$. As previously explained, we do not integrate $\bf b$} directly, but rather use $f({\bf b})={\bf A}^{\top}{\bf b}$ and embed this vector in the network.
@@ 2.5. Architectures for data-aware and noise-informed velocity estimators
be an encoder network that is parameterized by $\boldsymbol{\eta}$}. The choice of using $f({\bf b})={\bf A}^{\top}{\bf b}$ as the data embedding was primarily motivated by the need to map the measurement data back into the image domain, where the velocity estimation network operates. In our framework, the neural network is designed to predict the velocity field $s_{\boldsymbol{\theta}}({\bf x}_{t},f({\bf b}),t,\sigma)$ within the image space of $\bf x$}, whereas the measured data $\bf b$} (e.g., sinograms in tomography) often resides in a different domain with a different dimensionality. Also, this design makes the network more robust and better conditioned, because the embedding ${\bf A}^{\top}{\bf b}$ acts as a non-local feature map. For the tomography problem, each pixel receives contributions from multiple rays (or projections), effectively aggregating contextual information from the entire measurement domain. This non-local property allows the network to exploit long-range dependencies and global structures in the reconstruction, rather than relying solely on local features.
Formally, the network estimates the velocity as,
$\hat{\bf v}={\bf s}_{{\boldsymbol{\theta}}}({\bf x}_{t},E_{{\boldsymbol{\eta}}}({\bf A}^{\top}{\bf b}),t,\sigma).$ (16)
@@
2.6. Inference and uncertainty estimation
Figure 5. Computing uncertainty in the solutions obtained by FM for some example images from MNIST (left), STL10 (middle) and CIFAR10 (right) for the deblurring task. The reconstruction (posterior mean) is computed by averaging over runs from 32 randomly initialized ${\bf x}_{0}$. The uncertainty in the posterior mean is computed by evaluating its standard deviation for the predictions from 32 runs.
Our method is specifically aimed at highly ill-posed inverse problems, where regularization is often necessary to arrive at a stable solution. These problems often do not have a unique solution, and slight changes in input data can lead to large variations in the solution. Due to its stochastic nature, our method allows for the realization of multiple solutions since one can start with many random initial points ${\bf x}_{0}$ and evolve to multiple versions ${\bf x}_{1}$ of the solutions. This property can be used to generate even better estimates and for the estimation of uncertainty in the recovered images. To this end, assume that the ODE in Equation 10 is solved $M$ times starting each time at a different initial condition ${\bf x}_{0}^{(j)}$ and let
@@ 2.6. Inference and uncertainty estimation
${\bar{\bf x}}_{1}={\frac{1}{M}}\sum_{j=1}^{M}{\bf x}_{1}^{(j)},\quad{\sigma}_{{\bar{\bf x}}_{1}}={\frac{1}{M}}\sum_{j=1}^{M}\|{\bf x}_{1}^{(j)}-{\bar{\bf x}}_{1}\|^{2},$ (18)
as the mean and standard deviation of the estimated solutions. In particular, ${\bar{\bf x}}_{1}$ approximates the posterior mean and ${\sigma}_{{\bar{\bf x}}_{1}}$ estimates its standard deviation. It is well known that, for many problems the posterior mean can have a lower risk compared to other estimators (e.g. the Maximum A-Posteriori estimator). For an elaborate discussion on the properties of such estimators, see [Tenorio2011, somersallo, CalvettiSomersalo2005]. While such estimators are typically avoided due to computational complexity, the computational framework presented here allows us to compute them relatively easily.
@@
3. Details about the Inverse Problems
3.1. Details for image deblurring
$K(x,y)=\frac{1}{2\pi\sigma_{x}\sigma_{y}}\exp\bigg(-\frac{x^{2}}{\sigma_{x}^{2}}-\frac{y^{2}}{\sigma_{y}^{2}}\bigg).$ (20)
The convolution operation can be described as a multiplication in the Fourier domain,
$\tilde{B}(u,v)=\tilde{I}(u,v)\cdot\tilde{K}(u,v),$ (21)
where $\tilde{I}(u,v)=\mathcal{F}\{I(x,y)\}$, $\tilde{K}(u,v)=\mathcal{F}\{K(x,y)\}$ are the Fourier transforms of the sharp image, and the PSF, respectively and $\mathcal{F}$ represents the Fourier transform. The blurred image can then be obtained using the inverse Fourier transform,
$I_{\text{blur}}(x,y)=\mathcal{F}^{-1}\{\tilde{B}(u,v)\}.$ (22)
Moreover, $\bf A$$^{\top}$, the adjoint of operator $\bf A$}, which is used to compute the data-embedding ${\bf A}^{\top}{\bf b}$ for the network, for some data $\bf b$} is easy to obtain. Since the image blurring operator is symmetric and self-adjoint, we have ${\bf A}={\bf A}^{\top}$. In this work, we compare our method to diffusion-based method [chung2022diffusion] and InverseUNetODE [eliasof2024completedeeplearningmethod] to highlight the robustness of our method against competing baselines (see Section 4).
3.2. Details for tomography
In tomography, the goal is to reconstruct the internal image of an object from a series of projections (sinograms) taken at various angles. Mathematically, this problem is formulated as solving an inverse problem: the original image $I$ is mapped to its projections $S$ through the Radon transform $\bf A$}, and reconstruction consists of estimating $I$ from these measured projections.
In this work, we consider the standard parallel-beam tomography model, in which the forward problem can be described as,
$\text{vec}({S})={\bf A}\text{vec}(I_{\text{pad}})+\epsilon,$ (23)
where, $\text{vec}(\cdot)$ represents vectorization (flattening), $\text{vec}(S)\in\mathbb{R}^{N_{\text{angles}}\cdot N_{\text{detectors}}}$ is the flattened sinogram (the set of projections), $\text{vec}(I_{\text{pad}})\in\mathbb{R}^{N_{\text{pixels}}}$ is the padded and flattened image of dimension $N_{\text{pixels}}$, ${\bf A}\in\mathbb{R}^{(N_{\text{angles}}\cdot N_{\text{detectors}})\times N_{\text{pixels}}}$ is the tomography projection matrix which describes the interaction of each ray with every pixel, and $N_{\text{angles}}$, $N_{\text{detectors}}$, and $N_{\text{pixels}}$ represent the number of tomographic projection angles, detectors, and pixels (in the padded image), respectively. The 2D sinogram $S\in\mathbb{R}^{N_{\text{angles}}\times N_{\text{detectors}}}$ is then obtained by reshaping $\text{vec}(S)$. The term $\epsilon$ is the additive noise to the data, which was chosen as $p\%$ of the range of values in ${\bf A}\text{vec}(I_{\text{pad}})$, with $p$ lying between 0% and 20% during training. For our experiments, we set $N_{\text{angles}}=180$, and $N_{\text{detectors}}=2s+1$, where $s\times s$ is the dimension of the original unpadded images.
A simple approach to reconstruct the image from measured sinograms is backprojection, which uses the adjoint operator $\bf A$$^{\top}$ of the forward projection matrix. The reconstructed image is then,
$\text{vec}(I_{\text{pad}})={\bf A}^{\top}\text{vec}(S).$ (24)
The resulting $\text{vec}(I_{\text{pad}})$ is reshaped and cropped back to size $s\times s$. However, because $\bf A$$^{\top}$ is not the analytic inverse of $\bf A$}, the reconstruction is typically blurred. This distinction is fundamental: for tomography, ${\bf A}^{\dagger}\neq{\bf A}^{\top}$, where $\bf A$$^{\dagger}$ represents the Moore-Penrose pseudoinverse of the Radon transform.
Filtered Backprojection (FBP): To address the blurring inherent in plain backprojection, classical tomography uses Filtered Backprojection (FBP) - an analytic inversion technique that restores high-frequency information lost during the projection process. FBP involves two main steps:
  • Filtering: Each 1D projection $S(\phi,z)$ (for angle $\phi$ and detector coordinate $z$) is Fourier transformed and multiplied by a frequency-domain filter $|\omega|$. Commonly used filters include ramp and Shepp-Logan filters. This step compensates for the attenuation of high frequencies in the measurement process.
  • Backprojection: The filtered projections are then inverse Fourier transformed and backprojected into the image space over all projection angles.
Mathematically, the reconstruction can be written as,
$I(x,y)=\int_{0}^{\pi}\bigg[\int_{-\infty}^{\infty}\tilde{S}(\phi,\omega)|\omega|e^{2\pi i\omega z}d\omega\bigg]d\phi,$ (25)
where $\tilde{S}(\phi,\omega)$ is the Fourier transform of $S(\theta,z)$ with respect to $z=x\text{cos}\phi+y\text{sin}\phi$ and $\omega$ represents the ramp filter. This additional filtering step is what distinguishes $\bf A$$^{\dagger}$ from $\bf A$$^{\top}$: while $\bf A$$^{\top}$ simply redistributes projection values back into the image domain, FBP explicitly corrects for the frequency imbalance to yield a sharper reconstruction.
For this task, we used ${\bf A}^{\top}{\bf b}$ instead of the analytic inverse to inject data $\bf b$} into the network because (i) $\bf A$$^{\top}$ provides a simple way to inject measurement information into the image domain, ensuring that the reconstructed estimate is consistent with the geometry of the data (to avoid training instabilities arising from different-sized multi-channel inputs), and (ii) our neural flow-matching approach learns to refine the backprojected image and compensate for the blur that would otherwise arise, effectively learning a data-driven analogue of the analytic filter step in FBP. In other words, our method treats ${\bf A}^{\top}\text{vec}(S)$ as a physically consistent but blurred prior, which the neural network then refines into a high-fidelity reconstruction. This hybrid approach leverages the interpretability and geometric fidelity of classical reconstruction while exploiting the representational power of deep generative models to achieve superior reconstruction quality. In this work, we compare our method against the FBP baseline with the ramp filter (see Section 4).
4. Numerical Experiments